OpenStudy (anonymous):

The base of a rectangular box measures 5 feet by 5 feet. If the volume is 125 cubic feet then the height is how many feet?

OpenStudy (anonymous):

im not sure sweetheart.... try wolframalpha.com see if that helps

OpenStudy (anonymous):

5

OpenStudy (anonymous):

how did you get 5?

OpenStudy (anonymous):

volume of cube = side^3 side *side * side= 125 5*5*side=125 so, side= 125/25=5